# Overcoming the Thermal-Noise Limit of Room-Temperature Microwave Measurements by “Cavity Pre-cooling” with a Low-Noise Amplifier —Application to Time-resolved Electron Paramagnetic Resonance

### v1

@@ A.3 NOISE ANALYSIS
The depth of mode cooling was also estimated by fitting a bi-exponential function to the increase in measured noise power upon switching out the cooling LNA, as shown in Fig. 4 in the main paper. The best-fit curve, along with a standard error band, is depicted in Fig. 15; this determined $\Delta P_{\text{exp-Cool}}=-3.5\pm 0.4$ dB, agreeing both with our simulated $\Delta P_{\text{sim}}$ and our analysis using the noise spectrum from tr-EPR data, namely $\Delta P_{\text{exp-EPR}}$ as shown in Fig. 6(b) of the main text. Our estimates are summarized in Table 2.
@@
Figure 15: Error analysis and fitting of the experimental warming-up signal as presented in Fig. 4 of the main paper. The grey area represents the period when the double-throw switch is still transitioning to connect the over-coupled port to an “open” load, disconnecting the ACL from the resonator; this takes around $t_{\text{switch}}=2$ $\mu$s. Each blue dot plots $\Delta P_{\text{exp}}=10\log_{10}(<V^{2}_{\text{cold}}>/<V^{2}_{\text{ambient}}>)$ in units of dB[48], where $<V^{2}_{\text{cold}}>$ is the mean-squared received noise voltage averaged over ten consecutive data points, spaced 100 ns apart, spanning one microsecond; $<V^{2}_{\text{ambient}}>$ is the equivalent mean-squared noise voltage under ambient conditions (based on an adjoining sample spanning several hundred microseconds). The solid red line, along with its surrounding salmon pink standard error band, was obtained by least-squares fitting a bi-exponential to the evolution of $\Delta P_{\text{exp}}$ excluding the initial switching period. The longer of the two decay constants was determined to be $\tau_{\text{warm up}}^{\text{exp.}}\approx 9\pm 1\mu$s.
Table 2: Simulated and experimentally-inferred depth of cooling
method: Simulation Direct monitoring EPR data analysis
@@
label: sim exp-Cool exp-EPR
Noise-Power Reduction ($\Delta$ P) in dB: 3.5 3.5 3.7
Standard Deviation of $\Delta$ P: $\pm$ 0.2 $\pm$ 0.4 $\pm$ 1.0

### v2

@@ III Noise Analysis
The depth of mode cooling was also estimated by fitting a bi-exponential function to the increase in measured noise power upon switching out the cooling LNA, as shown in Figs. 4(c) and 7 in the main paper. The best-fit curve, with its associated standard error band, is depicted in Fig. S9; this determines $\Delta P_{\text{exp-Cool}}=-3.5\pm 0.4$ dB, agreeing (within error bars) of both our model’s prediction $\Delta P_{\text{model}}$ and our analysis using the noise spectrum from tr-EPR data, namely $\Delta P_{\text{exp-EPR}}$ as shown in Fig. 8(b) of the main text. Our estimates are summarized in Table S2.
@@
Figure S9: Error analysis and fitting of the experimental warming-up signal as presented in Fig. 6 of the main paper. The grey area at shorts times (after switching off the cooling LNA) on the far left represents the period when the double-throw switch is still transitioning to connect the over-coupled port to an “open” load, disconnecting the ACL from the resonator; this takes around $t_{\text{switch}}=2$ $\mu$s. Each blue dot plots $\Delta P_{\text{exp}}=10\log_{10}(<V^{2}_{\text{cold}}>/<V^{2}_{\text{ambient}}>)$ in units of dBSung et al. (2003), where $<V^{2}_{\text{cold}}>$ is the mean-squared received noise voltage averaged over ten consecutive data points, spaced 100 ns apart, spanning one microsecond; $<V^{2}_{\text{ambient}}>$ is the equivalent mean-squared noise voltage under ambient conditions (based on an adjoining sample spanning several hundred microseconds). The solid red line, along with its surrounding salmon pink standard error band, was obtained by least-squares fitting a bi-exponential to the evolution of $\Delta P_{\text{exp}}$ excluding the initial switching period. The longer of the two decay constants was determined to be $\tau_{\text{warm up}}^{\text{exp.}}\approx 9\pm 1\mu$s.
Table S2: Modeled and experimentally-inferred depth of cooling
method: Model Direct monitoring EPR data analysis
@@
label: model exp-Cool exp-EPR
Noise-power reduction ($\Delta P$ ) in dB: 3.2 3.5 3.7
Standard deviation of $\Delta P$ in dB: $\pm$ 0.2 $\pm$ 0.4 $\pm$ 1.0
Inferred mode temperature ($T$mode) in K 131.5 122.6 -
Error in $T$mode in K {+7,-6} {+13, -12} -